Metropolitan Casting Services started the year with total assets of $130,000 and total liabilities of $45,000. The revenues and the expenses for the year amounted to $130,000 and $50,000, respectively. During the year, the company did not issue any common stock, but it distributed dividends of $60,000. Calculate the amount of increase or decrease in stockholders' equity for the year.

A. a $85,000 decrease

B. a $60,000 increase

C. a $105,000 increase

D. a $20,000 increase
Answer #1

Answer: D) a $20,000 Increase

Total Stockholder's Equity in the Start = Total assets - Liabilities = 130,000-45,000 = 85,000

Increase in Stockholders equity

= Revenue -Expenses -Dividend

= 130,000-50,000-60,000 = $20,000 Increase